# Shortcuts to adiabaticity for coherent atom transport in an adjustable family of two-dimensional optical lattices

### v1

@@ VI Results and discussion
In what follows, we quantify and discuss the efficiency of STA-based coherent single-atom transport (cf. Sec. IV above) in various optical lattices belonging to the lattice family under consideration. The main figure of merit used for this purpose is the transport fidelity, defined as
$\mathcal{F}(t_{f})=|\braket{\Psi_{\textrm{target}}|\Psi(t_{f})}|^{2}\>.$ (39)
@@
Its dependence on the transport time $t_{f}$ is determined by the overlap of the target state $|\Psi_{\textrm{target}}\rangle$ (the ground state of the displaced optical-lattice potential, computed using the FSOM combined with the ITE approach; see Appendix B) and the actual final (time-evolved) atomic state $|\Psi(t_{f})\rangle$ (evaluated using FSOM-based numerical propagation of the TDSE in the comoving frame; see Sec. V above and Appendix A).
Figure 5: Dependence of the atom-transport fidelity on the transport time $t_{\mathrm{f}}$ for different transport distances $d_{x}$ and potential depths using the STA scheme. The ratio of the lattice depths corresponds to that of a honeycomb lattice. Panels (a)-(c) show results for potential depths of $(U_{X},U_{\overline{X}},U_{Y})=(200,600,200)E_{R}$, $(400,1200,400)E_{R}$, and $(800,2400,800)E_{R}$, respectively.
Figure 6: Dependence of the atom-transport fidelity on the transport time $t_{\mathrm{f}}$ for different transport distances $d_{x}$ using the STA scheme. The ratio of the lattice depths corresponds to that of a square lattice. Panels (a)-(c) show results for potential depths of $(U_{X},U_{\overline{X}},U_{Y})=(0,200,200)E_{R}$, $(0,400,400)E_{R}$, and $(0,800,800)E_{R}$, respectively.
Figure 7: Dependence of the atom-transport fidelity on the transport time $t_{\mathrm{f}}$ for different transport distances $d_{x}$ using the STA scheme. The ratio of the lattice depths corresponds to that of a linear lattice (1D chains). Panels (a)-(c) show results for potential depths of $(U_{X},U_{\overline{X}},U_{Y})=(50,1000,200)E_{R}$, $(100,2000,400)E_{R}$, and $(200,4000,800)E_{R}$, respectively.
Figure 8: Dependence of the atom-transport fidelity on the transport time $t_{\mathrm{f}}$ for different transport distances $d_{x}$ using the STA scheme. The ratio of the lattice depths corresponds to that of a dimerized lattice. Panels (a)-(c) show results for potential depths of $(U_{X},U_{\overline{X}},U_{Y})=(200,400,200)E_{R}$, $(400,800,400)E_{R}$, and $(800,1600,800)E_{R}$, respectively.
@@
The obtained results for the transport fidelity in the optical lattices of the adjustable family under consideration are depicted in Figs. 5-8. One common denominator of all these results, irrespective of the chosen values of the system parameters (such as the lattice depth, transport distance, beam waists, etc.), is the collapse of transport – manifested by a precipitous decrease in the fidelity – for very short transport times $t_{f}$. This characteristic breakdown of transport originates from the fact that for atomic accelerations above a certain maximal value $|a_{\textrm{max}}|$ that is proportional to the lattice depth Hau the atom being transported cannot be confined by the optical-lattice potential. This happens as a consequence of the disappearance of the minima of the “tilted” lattice potential experienced by the atom in its comoving (non-inertial) reference frame; compared to the potential in the lab frame, the comoving-frame potential has an additional term linear in the transport coordinate Hau . This characteristic behavior for short transport times is a generic property of coherent atom transport in optical lattices, i.e. it is not special for the concrete family of lattices being considered here.
@@
The aforementioned breakdown of transport occurs when the maximal acceleration of an atom during transport - hereafter denoted by $|\tilde{a}_{\textrm{max}}|$ - becomes larger than $|a_{\textrm{max}}|$. It should be borne in mind that the actual value of $|\tilde{a}_{\textrm{max}}|$ depends on the chosen dynamical-lattice trajectory (i.e. the path of the potential minimum); yet, quite generally it holds that $|\tilde{a}_{\textrm{max}}|\propto t_{f}^{-2}$. It immediately follows from this last result that for deeper lattices – i.e., for a larger value of $|a_{\textrm{max}}|$, which is proportional to the lattice depth – the largest atomic acceleration $|\tilde{a}_{\textrm{max}}|$ reached becomes larger than $|a_{\textrm{max}}|$ for sufficiently short total time $t_{f}$. In other words, for deeper lattices the breakdown of transport (i.e., the abrupt decrease of the transport fidelity) takes place for shorter times $t_{f}$.
Our obtained numerical results for the transport fidelity [cf. Eq. (39)] are displayed in Figs. 5-8 for honeycomb, square, linear (1D chain), and dimerized lattices, respectively, and two different transport distances ($d=100\>l_{x}$ and $d=400\>l_{x}$). These results are consistent with the above general features of coherent atom transport in optical lattices. In other words, they clearly manifest breakdown of transport at short transport times. On the other hand, in the opposite limit of very long transport times $t_{\mathrm{f}}$ the results coincide with the well-known adiabatic limit, i.e. essentially perfect atom transport ($\mathcal{F}\approx 1$) for $t_{\mathrm{f}}\gg\omega_{\mathrm{j}}^{-1}$ ($j=x,y$) is achieved.
In particular, in the case of the honeycomb lattice, with the parameter values chosen as in Fig. 5, the respective transport times $t_{f}$ corresponding to the onset of the breakdown of atom transport are in the range between $7\>T_{\mathrm{x}}$ and $8\>T_{\mathrm{x}}$ for the transport distance $d=100\>l_{x}$, as well as in the range between $16\>T_{\mathrm{x}}$ and $17.5\>T_{\mathrm{x}}$ for the transport distance $d=400\>l_{x}$.
In the case of the square lattice, which corresponds to Fig. 6, the breakdown of atom transport takes place at times in the range between $5.5\>T_{\mathrm{x}}$ and $6\>T_{\mathrm{x}}$ for the transport distance $d=100\>l_{x}$, and between $10\>T_{\mathrm{x}}$ and $12\>T_{\mathrm{x}}$ for $d=400\>l_{x}$. These results are clearly in agreement with the anticipated trend of decreasing with the increase of the lattice depth.
For the linear (1D-chain) lattice that corresponds to Fig. 7, the transport breakdown takes place when $t_{f}$ is between $5.5\>T_{\mathrm{x}}$ and $6\>T_{\mathrm{x}}$ for $d=100\>l_{x}$, and between $10\>T_{\mathrm{x}}$ and $12\>T_{\mathrm{x}}$ for $d=400\>l_{x}$.
Finally, in the case of the dimerized lattice with the parameter values chosen as in Fig. 8, the onset of the transport breakdown occurs for $t_{f}$ between $10\>T_{\mathrm{x}}$ and $13\>T_{\mathrm{x}}$ for the transport distance $d=100\>l_{x}$, and between $14\>T_{\mathrm{x}}$ and $18\>T_{\mathrm{x}}$ for $d=400\>l_{x}$.

### v2

@@ VI Results and discussion
In what follows, we quantify and discuss the efficiency of STA- and eSTA-based coherent single-atom transport (cf. Secs. IV.1 and IV.2 above, respectively) in various optical lattices belonging to the lattice family under consideration. The main figure of merit used for this purpose is the transport fidelity, defined as $\mathcal{F}(t_{f})=|\braket{\Psi_{\textrm{target}}|\Psi(t_{f})}|^{2}$. Its dependence on the transport time $t_{f}$ is determined by the overlap of the target state $|\Psi_{\textrm{target}}\rangle$ (the ground state of the displaced optical-lattice potential, computed using the FSOM combined with the ITE approach; see Appendix D) and the actual final (time-evolved) atomic state $|\Psi(t_{f})\rangle$ (evaluated using FSOM-based numerical propagation of the TDSE in the comoving frame; see Sec. V above and Appendix C).
In Sec. VI.1 below, we present our results for atom-transport fidelities. In Sec. VI.2 we then discuss how these results are affected in the presence of small variations in lattice depths. Finally, in Sec. VI.3 we demonstrate the superiority of our adopted transport schemes compared to approaches based on ad-hoc velocity profiles.
VI.1 Atom-transport fidelities
@@
The obtained results for the transport fidelity in the optical lattices under consideration are depicted in Figs. 512. One common denominator of all these results, irrespective of the chosen values of the system parameters (such as the lattice depth, transport distance, beam waists, etc.), is the collapse of transport – manifested by a precipitous decrease in the fidelity – for very short transport times $t_{\mathrm{f}}$. This characteristic breakdown of transport originates from the fact that for atomic accelerations above a certain maximal value $|a_{\textrm{max}}|$, which is proportional to the lattice depth Hau , the atom being transported cannot be confined by the optical-lattice potential. This happens as a consequence of the disappearance of the minima of the “tilted” lattice potential experienced by the atom in its comoving (non-inertial) reference frame; compared to the potential in the lab frame, the comoving-frame potential has an additional term linear in the transport coordinate Hau . This characteristic behavior for short transport times is a generic property of coherent atom transport in optical lattices, i.e. it is not special for the concrete family of lattices being considered here.
Figure 5: (Color online) Dependence of the STA- and eSTA atom-transport fidelity on the transport time $t_{\mathrm{f}}$ for a transport distances of $d_{\mathrm{x}}=100\,l_{\mathrm{x}}$ and fixed potential depths. The ratio of the lattice depths corresponds to that of a square lattice [cf. Fig. 2(a)]. Panels (a)-(c) show results for potential depths of $(U_{X},U_{\overline{X}},U_{Y})=(0,200,200)E_{R}$, $(0,400,400)E_{R}$, and $(0,800,800)E_{R}$, respectively.
Figure 6: (Color online) Dependence of the STA- and eSTA atom-transport fidelity on the transport time $t_{\mathrm{f}}$ for a transport distances of $d_{\mathrm{y}}=100\,l_{\mathrm{y}}$ and fixed potential depths. The ratio of the lattice depths corresponds to that of a square lattice [cf. Fig. 2(a)]. Panels (a)-(c) show results for potential depths of $(U_{X},U_{\overline{X}},U_{Y})=(0,200,200)E_{R}$, $(0,400,400)E_{R}$, and $(0,800,800)E_{R}$, respectively.
@@
The collapse of transport actually takes place when the maximal acceleration of an atom during the transport process, to be hereafter denoted by $|\tilde{a}_{\textrm{max}}|$, exceeds $|a_{\textrm{max}}|$. Needless to say, the actual value of $|\tilde{a}_{\textrm{max}}|$ depends upon the chosen dynamical-lattice trajectory (i.e. the path of the potential minimum); however, it holds quite generally that $|\tilde{a}_{\textrm{max}}|\propto\>t_{\mathrm{f}}^{-2}$. What follows from this last result is that for deeper lattices, which then also implies a larger value of $|a_{\textrm{max}}|$ because the latter is proportional to the lattice depth, the largest atomic acceleration $|\tilde{a}_{\textrm{max}}|$ reached exceeds $|a_{\textrm{max}}|$ for sufficiently short transport times $t_{\mathrm{f}}$. In other words, for deeper lattices the collapse of transport (i.e., the precipitous decrease of the transport fidelity) takes place for shorter times $t_{\mathrm{f}}$. Our numerical results are consistent with this general argument, as can be verified for gradually increasing lattice depths in all figures independent on the specific transport direction or lattice configuration.
In the opposite limit of very long transport times $t_{\mathrm{f}}$, the STA results converge to the well-known adiabatic regime. In this limit, essentially perfect atom transport ($\mathcal{F}\approx 1$) can be achieved for $t_{\mathrm{f}}\gg\omega_{\mathrm{u}}^{-1}$ ($u=x,y$). However, this behavior is only partially recovered in the eSTA approach. In cases where the potential minimum does not reside at the origin—corresponding to the presence of an additional linear term in the harmonic approximation, cf. Eq. (12)—the adiabatic limit is reached only for sufficiently deep lattice potentials, as discussed for the honeycomb and dimer potentials at the end of this section.
The transport fidelity corresponding to the square lattice for displacements along the $x$-direction is depicted in Fig. 5. While the eSTA fidelity curve in this case exhibits a pronounced dip for shallow lattice potentials in panel (a), it nevertheless achieves a significantly improved fidelity compared to its STA counterpart for transport times $t_{\mathrm{f}}\approx 3.8\,T_{\mathrm{X}}$. This improvement becomes increasingly pronounced for deeper potentials, culminating in an almost uniformly high fidelity close to unity in panel (c). The remaining minor deviations are a shallow dip for $t_{\mathrm{f}}\approx 4.2\,T_{\mathrm{X}}$ and a slightly earlier onset of fidelity degradation for $t_{\mathrm{f}}\approx 3.2\,T_{\mathrm{X}}$. For even deeper lattices, these features are expected to vanish completely, yielding a consistently improved transport performance through the eSTA scheme for the square lattice. Owing to the high degree of symmetry of this lattice configuration [cf. Fig. 2(a)], an analogous behavior is observed for displacements along the $y$-direction, as illustrated in Fig. 6.
For displacements along the $y$-direction in the 1D-chains lattice the corresponding STA- and eSTA fidelity curves are shown in Fig. 8. What can be inferred from these results is that their qualitative behavior closely follows that of the square lattice. The most notable differences are the minimal achievable high-fidelity transport times, which are shifted to $t_{\mathrm{f}}\approx(6.5-7.5)\,T_{\mathrm{X}}$, and the increased lattice depth of $U_{\mathrm{\overline{X}}}$, reaching values up to $4000\,E_{R}$. While this depth is substantially larger than in the square-lattice case, it remains within reach of state-of-the-art experimental optical-lattice setups.
Figure 7: (Color online) Dependence of the STA- and eSTA atom-transport fidelity on the transport time $t_{\mathrm{f}}$ for a transport distances of $d_{\mathrm{x}}=100\,l_{\mathrm{x}}$ and fixed potential depths. The ratio of the lattice depths corresponds to that of the 1D-chains lattice [cf. Fig. 2(d)]. Panels (a)-(c) show results for potential depths of $(U_{X},U_{\overline{X}},U_{Y})=(50,1000,200)E_{R}$, $(100,2000,400)E_{R}$, and $(200,4000,800)E_{R}$, respectively.
Figure 8: (Color online) Dependence of the STA- and eSTA atom-transport fidelity on the transport time $t_{\mathrm{f}}$ for a transport distances of $d_{\mathrm{y}}=100\,l_{\mathrm{y}}$ and fixed potential depths. The ratio of the lattice depths corresponds to that of the 1D-chains lattice [cf. Fig. 2(d)]. Panels (a)-(c) show results for potential depths of $(U_{X},U_{\overline{X}},U_{Y})=(50,1000,200)E_{R}$, $(100,2000,400)E_{R}$, and $(200,4000,800)E_{R}$, respectively.
By contrast to atom transport in the $y$ direction, transport along the $x$-direction of the 1D-chains lattice [for an illustration, see Fig. 7, exhibits a qualitatively different behavior. Rather than displaying a sharp dip followed by an abrupt loss of fidelity, the STA transport fidelity is characterized by a more gradual degradation as the transport time $t_{\mathrm{f}}$ decreases. This can be attributed to the more slowly varying potential depth along the $x$-direction. Remarkably, the eSTA transport fidelity already provides a noticeable improvement for relatively shallow lattice depths in panel (a), albeit with a minor dip for $t_{\mathrm{f}}\approx 4.3\,T_{\mathrm{X}}$. For increasing lattice depths, eSTA consistently outperforms STA across the entire range of transport times, where even the previously described dip in fidelity is not visible anymore. It is important to notice that the sharp fidelity drop for the eSTA case observed in other geometries for small $t_{\mathrm{f}}$ is absent here; instead, eSTA transport fidelity exhibits a smoother decay as well, closely mirroring the gradual behavior of its STA counterpart.
The STA results for transport along the $x$-direction of the honeycomb lattice, shown in Fig. 9, similarly display a gradual reduction in transport fidelity for shorter transport times and do not exhibit a pronounced dip prior to the final drop. In addition, small oscillations are observed for $t_{\mathrm{f}}\approx 4.8\,T_{\mathrm{X}}$ and $t_{\mathrm{f}}\approx 5.5\,T_{\mathrm{X}}$ for all considered lattice depths, panels (a)–(c). These oscillations take place at times that are independent of the lattice depth, while their amplitude decreases for deeper potentials. These features can be attributed to the non-vanishing linear term in the harmonic approximation, Eq. (12), which induces minor instabilities in the eSTA-based transport scheme, particularly for shallower lattices.
Figure 9: (Color online) Dependence of the STA- and eSTA atom-transport fidelity on the transport time $t_{\mathrm{f}}$ for a transport distances of $d_{\mathrm{x}}=50\,l_{\mathrm{x}}$ and fixed potential depths. The ratio of the lattice depths corresponds to that of a honeycomb lattice [cf. Fig. 2(c)]. Panels (a)-(c) show results for potential depths of $(U_{X},U_{\overline{X}},U_{Y})=(400,1200,400)E_{R}$, $(800,2400,800)E_{R}$, and $(1600,4800,1600)E_{R}$, respectively.
Figure 10: (Color online) Dependence of the STA-and eSTA atom-transport fidelity on the transport time $t_{\mathrm{f}}$ for a transport distances of $d_{\mathrm{y}}=50\,l_{\mathrm{y}}$ and fixed potential depths. The ratio of the lattice depths corresponds to that of a honeycomb lattice [cf. Fig. 2(c)]. Panels (a)-(c) show results for potential depths of $(U_{X},U_{\overline{X}},U_{Y})=(400,1200,400)E_{R}$, $(800,2400,800)E_{R}$, and $(1600,4800,1600)E_{R}$, respectively.
A similar discussion applies to transport along the $y$-direction of the honeycomb lattice, illustrated by the transport-fidelity curves in Fig. 10. Although the fidelity drop at short transport times occurs more abruptly and at slightly smaller values of $t_{\mathrm{f}}$ [more precisely, for $t_{\mathrm{f}}\approx(2.2-2.8)\>T_{\mathrm{X}}$], the eSTA approach nevertheless outperforms the STA scheme in this regime.
Finally, transport in the dimer lattice along both the $x$- and $y$-directions, respectively illustrated by the fidelity curves in Figs. 11 and 12, remains unstable even for extremely deep lattice potentials. This behavior can be partially attributed to the enhanced contribution of the linear term in the harmonic approximation, or, equivalently, to the large values of $x_{0}$ for this potential configuration [cf. Eq. (7)]. These effects give rise to oscillations that are responsible for the fidelity dips observed for $t_{\mathrm{f}}\approx 6\,T_{\mathrm{X}}$ and $t_{\mathrm{f}}\approx 7.1\,T_{\mathrm{X}}$. In addition, the exceptionally shallow potential barrier separating the two dimer sites plays an important role. As illustrated in Fig. 2(b), this barrier is extremely shallow and increases only marginally with increasing lattice depth. Although eSTA exhibits some improvement for deeper lattices, it does not surpass the STA performance in this case. We expect that eSTA would eventually outperform STA for even deeper potentials; however, computational limitations prevent us from accessing this regime even further.
Figure 11: (Color online) Dependence of the STA- and eSTA atom-transport fidelity on the transport time $t_{\mathrm{f}}$ for a transport distances of $d_{\mathrm{x}}=50\,l_{\mathrm{x}}$ and fixed potential depths. The ratio of the lattice depths corresponds to that of a dimer lattice [cf. Fig. 2(b)]. Panels (a)-(c) show results for potential depths of $(U_{X},U_{\overline{X}},U_{Y})=(12800,25600,12800)E_{R}$, $(25600,51200,25600)E_{R}$, and $(51200,102400,51200)E_{R}$, respectively.
Figure 12: (Color online) Dependence of the STA- and eSTA atom-transport fidelity on the transport time $t_{\mathrm{f}}$ for a transport distances of $d_{\mathrm{y}}=50\,l_{\mathrm{y}}$ and fixed potential depths. The ratio of the lattice depths corresponds to that of a dimer lattice [cf. Fig. 2(b)]. Panels (a)-(c) show results for potential depths of $(U_{X},U_{\overline{X}},U_{Y})=(12800,25600,12800)E_{R}$, $(25600,51200,25600)E_{R}$, and $(51200,102400,51200)E_{R}$, respectively.
These effects are most pronounced for eSTA-based transport along the $x$-direction, while being noticeably improved – but not entirely suppressed – for displacements along $y$, as demonstrated by the eSTA results in Fig. 12. Thus, although transport in the $y$-direction exhibits enhanced stability compared to the $x$-direction, residual instabilities remain and do not vanish completely. By contrast, STA-based transport does not exhibit instabilities for the dimer lattice in either the $x$- or $y$-direction.
It should be emphasized that the eSTA results for the dimer lattice require unrealistically deep potentials that are far beyond current experimental capabilities. Consequently, it serves primarily as an academic test case to elucidate the challenging regimes of the eSTA protocol. For this reason, the dimer lattice will not be considered further in the subsequent sections, which focus on experimentally relevant configurations.
Figure 13: (Color online) Relative stability to reach an atom-transport fidelity of $0.99$, defined as $1-t_{\mathrm{real}}/t_{\mathrm{id.}}$, as a function of the instability parameter $\xi$. Here, $t_{0.99}$ is the transport time that first drops below $0.99$ for a given $\xi$ while $t_{\mathrm{id.}}$ is the reference time $t_{0.99}$ for $\xi=0$. Curves are shown for both the STA and eSTA schemes, illustrating how the transport time is affected by lattice instabilities. The lattice depth ratios correspond to (a) the square lattice, with $(U_{X},U_{\overline{X}},U_{Y})=(0,800,800)\,E_{R}$; (b) the 1D-chains lattice, with $(U_{X},U_{\overline{X}},U_{Y})=(100,2000,400)\,E_{R}$; (c) the honeycomb lattice $(U_{X},U_{\overline{X}},U_{Y})=(1600,4800,1600)\,E_{R}$. The transport distance is $d_{\mathrm{x}}=100\,l_{\mathrm{x}}$ for the square and 1D-chains lattice, and $d_{\mathrm{x}}=50\,l_{\mathrm{x}}$ for the honeycomb lattice.
VI.2 Robustness against deviations in lattice depths
To account for possible imperfections in envisioned experimental implementations of our proposed transport schemes, here we address the sensitivity of those schemes to deviations in lattice depths. It is pertinent to assume that in an actual implementation of our STA- and eSTA transport schemes the nominal potential depths $U_{j}$ ($j\in\{X,\overline{X},Y\}$) take slightly changed values $\widetilde{U}_{j}=U_{j}\left(1+\xi\right)$, where $\xi$ denotes the instability parameter.
To simplify our analysis, we keep $\widetilde{U}_{j}$ fixed at some $\xi$-independent values and make use of those values as the actual potential depths in our evaluation of STA- and eSTA transport fidelities; we previously design STA- and eSTA dynamical-lattice trajectories (recall Sec. IV) with $U_{j}(\xi)\equiv\widetilde{U}_{j}/(1+\xi)$ playing the role of potential depths for each of the considered values of $\xi$. The advantage of this approach is that by fixing the potential depth the maximal acceleration supported by the potential remains the same for all data points (i.e. for different values of $\xi$), which facilitates a clearer assessment of sensitivity to deviations. While in the previous section a direct, one-to-one comparison between STA and eSTA schemes was central, here the focus lies on the individual stability properties of each scheme under parameter variations.
The results of our analysis are illustrated in Fig. 13, where the relative deviation $1-t_{0.99}/t_{\mathrm{id.}}$ is plotted as a function of the instability parameter $\xi$. Here, $t_{0.99}$ denotes the transport time at which the fidelity first drops below $0.99$, while the reference time $t_{\mathrm{id.}}$ corresponds to the respective scheme (STA or eSTA) evaluated at $\xi=0$. By construction, all curves pass through the origin.
A notable trend in the STA results, visible across all three panels, is that increasing the instability parameter leads to a reduced transport time $t_{0.99}$, resulting in an approximately constant negative slope. This behavior can be attributed to the fact that decreasing $\xi$ effectively reduces the lattice depths and, consequently, the frequencies of the harmonic approximations underlying the STA construction. At shorter transport times $t_{\mathrm{f}}$, where higher excited states contribute more strongly, a lower effective frequency provides a better approximation of the potential far away from its minimum. In this sense, reducing the frequency can be interpreted as a heuristic improvement that captures the overall potential shape more faithfully than a harmonic approximation that is better suited around the origin.
An exception to the aforementioned trend is found in the honeycomb-lattice case shown in panel (c), where a sudden increase in $t_{0.99}$ occurs around $\xi\approx-0.4$. This behavior likely originates from the fact that the harmonic frequencies are becoming too large to reliably represent the global structure of the underlying potential.
Overall, the eSTA results follow a similar qualitative trend as the STA scheme, which is expected given that eSTA is constructed as a potential tailored version build on STA. Nevertheless, additional features emerge. For the square lattice in panel (a), a pronounced reduction of $t_{0.99}$ is observed near $\xi=0$, extending into the regime of negative instability parameters till around 0.2-0.2. In this interval, the eSTA protocol consistently outperforms the reference time $t_{\mathrm{id.}}$, before reverting to the general trend. Since the eSTA scheme is more strongly tailored to the reference potential, and is thus expected to be more sensitive to deviations, this region of enhanced stability does not extend over the entirety of the probed regime.
Similar behavior is observed for the remaining lattice geometries. Notably, for the 1D-chains lattice shown in panel (b), the point $\xi=0$ itself lies within the region of improved stability. This indicates that incorporating small deviations from the reference potential in the construction of the control protocol can enhance robustness, although the extent and location of this improvement depend sensitively on the specific lattice geometry.
VI.3 Comparison with atom transport based on ad-hoc velocity profiles
Figure 14: (Color online) Dependence of the atom-transport fidelity on the transport time $t_{\mathrm{f}}$ for fixed potential depths using the STA- and eSTA-based transport, as well as transport based on sine- and triangular (ad-hoc) velocity profiles. The ratio of the lattice depths corresponds to that (a) the square lattice with $(U_{X},U_{\overline{X}},U_{Y})=(0,800,800)E_{R}$; (b) the 1D-chains lattice with $(U_{X},U_{\overline{X}},U_{Y})=(200,4000,800)E_{R}$; (c) the honeycomb lattice with $(U_{X},U_{\overline{X}},U_{Y})=(1600,4800,1600)E_{R}$. The transport distance is $d_{\mathrm{x}}=100\,l_{\mathrm{x}}$ for the square and 1D-chains lattice, and $d_{\mathrm{x}}=50\,l_{\mathrm{x}}$ for the honeycomb lattice.
In the following, we conclude our analysis of STA and eSTA-based atom transport by comparing the results for the atom-transport fidelity obtained using these approaches with those corresponding to ad-hoc (independent of the optical-lattice potential) velocity profiles.
More specifically yet, we make comparisons between our STA- and eSTA results for atom transport to those based on the sine-shaped and triangular velocity profiles. The respective time-dependent paths $q_{0}^{\mathrm{triang}}(t)$ and $q_{0}^{\mathrm{sin}}(t)$ of the potential minimum corresponding to the triangular and sine-shaped velocity profiles are given by
$q_{0}^{\mathrm{triang}}(t)$ $=\begin{cases}v_{0}\,t^{2}/t_{f}\>,&\;\;0\leq t\leq t_{f}/2\\ v_{0}\left(2\,t-t_{f}/2-t^{2}/t_{f}\right)\>,&\;\;t_{f}/2<t\leq t_{f}\end{cases}\>,$ (53)
$q_{0}^{\mathrm{sine}}(t)$ $=\frac{v_{0}}{2}\left[t-\frac{\sin\left(2\pi\frac{t}{t_{f}}\right)}{2\pi}\>t_{f}\right]\>,$ (54)
where $v_{0}=2d/t_{f}$ is the maximal velocity during transport over the distance $d$.
The atom-transport approach based on the triangular velocity profile [cf. Eq. (53)] is often referred to as the bang-bang approach Tor ; Che (b); Din . This approach is accompanied by significant motional-heating effects in the fast-transport regime, a phenomenon inextricably linked with the presence of discontinuities in its corresponding acceleration profile. Consequently, such an approach leads to relatively inefficient atom transport in most practically relevant cases Hic . Somewhat better results than the ones originating from the bang-bang approach are obtained for atom transport based on the sine-shaped velocity profile [cf. Eq. (54)]; this is intimately related to the fact that its attendant acceleration of an atom remains continuous throughout the transport process.
By contrast to atom-transport approaches based on ad-hoc velocity profiles, the inverse-engineering based STA approach, and its eSTA counterpart to an even larger extent, make use of the specific form of the optical-lattice potential in question to obtain a tailored dynamical-lattice trajectory. For this reason, it is plausible to expect that these last two approaches should lead to much more time-efficient and robust atom transport than those based on pre-selected velocity profiles.
For the optical-lattice family under consideration, a comparison between transport based on ad-hoc velocity profiles and transport based on STA- and eSTA approaches is illustrated in Fig. 14 for the cases of the square-, 1D-chains-, and honeycomb lattice. As is evident from Fig. 14, the obtained results for the transport fidelity in the case of the triangular velocity show strong oscillations. The results obtained in the case of the sine-shaped velocity profile are closer – but still inferior – to those originating from the STA- and eSTA approaches. From the results obtained using the latter approaches it can be inferred that only one major drop in fidelity takes place in the case of STA-based transport in the square lattice [cf. Fig. 14(a)], while for the remaining two lattice configurations [cf. Fig. 14(b) and (c)] such drops are completely absent for both STA and eSTA transport fidelities.